# FLARE: Fast Low-rank Attention Routing Engine

### v2

@@ G.1 Introduction & Background
$\underbrace{\rho C_{p}\frac{\mathrm{d}T}{\mathrm{d}t}=\gradient\cdot k\Delta T(\bm{x},t)+Q(\bm{x},t),}_{\text{thermal transport}}\hskip 20.00003pt\underbrace{\gradient\cdot\bm{\sigma}=0{\color[rgb]{1,1,1}\frac{1}{2}}\hskip 10.00002pt\bm{\sigma}=\bm{C}\bm{\varepsilon}_{e},}_{\text{stress equilibrium}}$ (25)
@@ G.3 Benchmark task
G.4 Data filtering
The dataset contains a wide-varying range of shapes, making the data set general enough to train a strong data-driven field prediction model. Out of $\sim$27,000 shapes, 19,732 were successfully simulated. We analyze the first 3,500 successful simulations and filter them according to several statistics to design a balanced training and test set. For example, we limit the learning problem to meshes with up to 50,000 points and up to 300,000 edges. This is done to reduce memory usage which becomes a bottleneck when training on small GPUs. We also filter meshes that have high aspect ratio elements as the FEM calculation could be unreliable on highly distorted geometries. Our dataset processing code is available in the GitHub repository associated with this paper. The statistics for the filtered data set are presented in Table 6 along with histograms in Figure 12. A gallery of successful simulations is presented in Figure 11.

### v3

@@ H.1 Introduction & Background
$\underbrace{\rho C_{p}\frac{\mathrm{d}T}{\mathrm{d}t}=\gradient\cdot k\Delta T(\bm{x},t)+Q(\bm{x},t),}_{\text{thermal transport}}\hskip 20.00003pt\underbrace{\gradient\cdot\bm{\sigma}=0{\color[rgb]{1,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,1,1}\pgfsys@color@gray@stroke{1}\pgfsys@color@gray@fill{1}\frac{1}{2}}\hskip 10.00002pt\bm{\sigma}=\bm{C}\bm{\varepsilon}_{e},}_{\text{stress equilibrium}}$ (25)
@@
Figure 16: Qualitative results for FLARE on the LPBF $Z$-displacement prediction task. For each test geometry, we show the ground truth $Z$-displacement field, the model prediction, and the corresponding error (Ground Truth - Prediction).
H.2 Dataset Generation
H.3 Benchmark task
H.4 Data filtering
The dataset contains a wide-ranging set of shapes, making the dataset general enough to train a strong data-driven field prediction model. Out of $\sim$27,000 shapes, 19,732 were successfully simulated. We analyze the first 3,500 successful simulations and filter them according to several statistics to design a balanced training and test set. For example, we limit the learning problem to meshes with up to 50,000 points and up to 300,000 edges. This is done to reduce memory usage which becomes a bottleneck when training on small GPUs. We also filter meshes that have high aspect ratio elements as the FEM calculation could be unreliable on highly distorted geometries. Our dataset processing code is available in the GitHub repository associated with this paper. The statistics for the filtered dataset are presented in Table 6 along with histograms in Figure 15. A gallery of successful simulations is presented in Figure 14.
H.5 Qualitative Results for LPBF $Z$-Displacement Prediction
In Figure 16, we present visualizations of ground-truth $Z$-displacement, predictions by FLARE, and the corresponding error. Across a set of representative geometries from the LPBF test set, FLARE produces displacement fields that are visually indistinguishable from the ground truth and capture both the global deformation patterns and localized high-gradient regions. Importantly, the error fields remain low-magnitude and spatially diffuse, with no systematic accumulation near edges or corners—indicating that the model does not rely on positional shortcuts or overfit to particular geometric artifacts. This suggests that FLARE successfully learns the dominant thermo-mechanical modes of deformation while only missing small higher-order variations that are difficult to resolve without substantially larger latent bottlenecks. These qualitative observations are consistent with the low average relative $L^{2}$ error and illustrate that FLARE maintains stable accuracy across a diverse distribution of part geometries.
H.6 Dataset split and benchmark protocol
Splits.
From the full set of 19,732 successful LPBF simulations, we construct a filtered subset of cases with up to 50,000 points per geometry. We use 1,100 cases for training and 290 cases for testing, as summarized in Table 3.
Inputs and outputs.
For each case, the benchmark input consists of the 3D mesh node coordinates $\in\mathbb{R}^{N\times 3}$. The benchmark target is the final-time-step vertical displacement field ($Z$-displacement) $\in\mathbb{R}^{N\times 1}$ at the same nodes. We focus on this scalar displacement because it is directly tied to recoater-blade interference risk in LPBF.
Simulation context.
All parts are scaled to lie within a standardized build volume $30,30]\times[-30,30]\times[0,60]~\text{mm}$ and simulated in Autodesk NetFabb for the Renishaw AM250 machine and Ti-6Al-4V material system, using $40~\mu\text{m}$ layer thickness and layer lumping with a lumped layer height of $2.5~\text{mm}$. These settings follow the procedure described in Appendix H.